A farmer is trying to develop a feed mixture consisting of hay and grain. how many tons of hay at $110 per ton and how many tons of grain at $60 must be mixed to produce 100 tons of feed to sell at $70 per ton?
Answer by lwsshak3(11628)

let x = tons of grain at $60/ton
(100-x)= tons of hay at $110/ton
mixing grain and hay to produce 100 tons at $70/ton
=60x+110(100-x)=70(100)=7000
=60x+11000-110x=7000
50x=4000
x=4000/50=80
100-x=100-80=20
ans: 80 tons of grain at $60/ton mixed with 20 tons of hay at $110/ton would sell 100 tons of feed at $70/ton